# Tiered Reward Functions: Specifying and Fast Learning of Desired Behavior

### v2

@@ Abstract
Reinforcement-learning agents seek to maximize a reward signal through environmental interactions. As humans, our job in the learning process is to design reward functions to express desired behavior and enable the agent to learn such behavior swiftly. In this work, we consider the reward-design problem in tasks formulated as reaching desirable states and avoiding undesirable states. To start, we propose a strict partial ordering of the policy space to resolve trade-offs in behavior preference. We prefer policies that reach the good states faster and with higher probability while avoiding the bad states longer. Next, we introduce Tiered Reward, a class of environment-independent reward functions and show it is guaranteed to induce policies that are Pareto-optimal according to our preference relation. Finally, we demonstrate that Tiered Reward can lead to fast learning by evaluating on several environments using multiple tabular and deep reinforcement-learning algorithms.
@@ 1 Introduction
Reinforcement learning (Sutton and Barto 1998), aka RL, is concerned with the problem of learning to behave to maximize a reward signal. In biological systems, this reward signal is considered to be the organism’s motivational system, using pain and pleasure to modulate behavior. In engineered systems, however, rewards must be selected by the system designer. We view rewards as a kind of programming language—a specification of the agent’s target behavior (Littman et al. 2017). As arbiters of behavior correctness in the learning process, humans bear the responsibility of authoring this program.
@@
There are two essential steps in designing reward functions. First, one must decide what kind of behavior is desirable and should be conveyed. Then, there’s the choice of reward function that induces such behavior.
In this paper, we look at a specification language that allows for the expression of desirable states (goals and subgoals) and undesirable states (obstacles). Even in this simple setting, providing precise trade-offs is difficult. Is it better for an agent to increase the chance of getting to the goal by 5% if it also incurs 8% higher probability of hitting an obstacle? Is it better to increase by 50% the probability of getting to a goal if the expected time of getting there also increases by 20%? There is no universal preference over behavior and having to explicitly write down all possible trade-offs is challenging. Even if the reward designer has a way of expressing preferences for all possible exchanges, it can be difficult, or even impossible, to design a reward function that captures them without prior knowledge of the environment.
Favorable Policies
In the goal–obstacle class of tasks we consider, preferences over policies are simplest in the deterministic setting. We imagine all states are either goal states, obstacle states, or neither (background states). All goal states and obstacle states are absorbing. Preferences in deterministic environments form a total order: reaching the goal is better than reaching the obstacle; if the policy reaches the goal, faster is better; if the policy reaches the obstacle, avoiding it longer is better.
@@
As pointed out before, preferences are less clear in a stochastic setting because there can be trade-offs between different outcomes and their probabilities. However, some comparisons are arguably clear cut. Informally, if one policy induces uniformly better outcomes than another—being more likely to reach a goal and doing so faster, being less likely to reach an obstacle and getting there more slowly—we prefer such a policy. If the policies can’t be directly compared, we propose to be indifferent between them. Thus, we replace the standard reinforcement-learning notion of optimality with Pareto-optimality (Mornati 2013)—seeking a policy that is either preferred or incomparable to every other policy. Pareto-optimal policies are commonly adopted in the subfield of multiobjective RL (Vamplew et al. 2011).
Reward Design
@@
Policies in general are hard to express through reward functions (Amodei and Clark 2016), and some are even impossible to convey with a Markov (state–action-based) reward (Abel et al. 2021). Even when policies are expressible, designing bad reward functions can lead to undesirable or dangerous actions (Amodei and Clark 2016), easy reward hacking (Amodei et al. 2016), and more. We seek to design good reward functions, which can be characterized by many properties, such as interpretability and learning speed (Devidze et al. 2021). But the most important property a reward function must have is to guarantee the adoption of a desired policy. As we will show later in Section 5, even intuitively correct reward designs can lead to suboptimal policies. To hedge against this, we introduce a tiered reward structure that is guaranteed to induce Pareto-optimal policies. Intuitively, we partition the state space into several tiers, or goodness levels. States in the same tier are associated with the same reward, while states in a more desirable tier are associated with a proportionally higher reward. We prove that these tiered reward structures, with the proper constraints between reward values, induce Pareto-optimal behavior and empirically show that they can lead to fast learning.
@@
Our contribution is threefold: First, we define a preference over the entire policy space via a strict partial ordering on outcomes. Then, we introduce a class of environment-independent tiered reward structure that provably induce Pareto-optimal policies with respect to this preference ordering. Finally, we demonstrate these tiered reward functions can lead to fast learning in both tabular and deep RL settings.
@@ 2 Related Work
There are many papers that deal with specifying behavior through rewards. Reward machines (Icarte et al. 2018, 2022) are finite state machines that compose reward functions and allow different rewards to be delivered dependent on the agent’s trajectory. They reveal the structure of the reward function to the RL agent to support decomposition of complex tasks. Our focus on how to provide incentives for specific outcomes is complementary and the two approaches can be used in concert. Temporal logic based languages (Littman et al. 2017; Camacho et al. 2017; Li, Vasile, and Belta 2017; Camacho et al. 2019) have been used to specify behavior. Though these methods can be more expressive, they often lead to intractable planning and learning problems due to state-space explosion issues (Wongpiromsarn, Topcu, and Murray 2010). We offer a different expressibility–tractability tradeoff. Preference-based RL methods (Wirth et al. 2017; Brown et al. 2019) learn a reward function based on a dataset of preferences over trajectories. But, as we have argued, preferences over trajectory probabilities can be very difficult to specify. In addition, our reward scheme relieves the need for environment-specific preference datasets created by human experts. Multi-objective RL (Vamplew et al. 2011; Toro Icarte et al. 2018) allows for different tasks to be specified through a set of reward functions. Our work proceeds in the orthogonal direction by designing a single reward function to trade off among multiple behaviors, instead of incentivizing all of them.
Designing reward functions to increase the speed of learning is a relatively open question. Mataric (1994) proposed to accelerate learning by incorporating domain knowledge and using a progress estimator, but does not provide a principled method of designing a reward function. Sowerby, Zhou, and Littman (2022) showed that reward functions that maximize the action gap given a measure of horizon length lead to fast learning. However, designing such reward functions require detailed knowledge of the environment, making this approach impractical for new learning problems.
@@ 3 Problem Setting
We are interested in the reward-design problem. In the common RL framework, tasks are specified by the reward function and the agent’s objective is to maximize rewards. We take an alternative perspective: We specify tasks by prescribing a set of desirable policies and seek reward functions such that maximizing the reward will lead to the desirable policies. Formally, given a set of desirable (Pareto-optimal) policies $\Pi$, the reward-design problem is to create a reward function $R:S\rightarrow R$ such that the optimal policy $\pi_{R}^{*}\in\Pi$.
We imagine the state space $S$ as exhibiting a tiered structure, where higher tiers are more desirable than lower tiers, and states within the same tier are equally desirable. We formally define a Tier MDP as:
@@
Figure 1: Russell/Norvig grid world. Objective is to reach the goal (green) without first visiting lava (red) ($\gamma=0.9$).
@@
As an example, the grid world from Russell and Norvig (2010), as illustrated in Figure 1, could be formulated as a 3-Tier MDP—the goal state is one tier ($S_{3}$), the lava state one tier ($S_{1}$), and all other states reside in the background tier ($S_{2}$). It is important to note that we put no constraints on how many states could be in each tier, nor how many tiers there can be. Therefore, the framework has a high degree of generality; any finite MDP with reward defined on states could be formulated as a Tier MDP by placing states with the same reward in the same tier. However, the Tier MDP is most useful when there are clear good and bad states in the state space, such as when there are goal and obstacle states, or even states of intermediate desirability such as subgoal states. In the following sections, we will show how to perform reward design in Tier MDPs.
4 Policy Ordering
@@
A policy can be thought of as inducing a probability distribution over an infinite set of outcomes (specifically the probability of reaching each of the states after $t$ steps, for all $t$). In goal–obstacle tasks, policies can be characterized by statistics such as probability of reaching the goal and probability of avoiding the obstacle for each possible horizon length.
For the moment, we limit the problem space to 3-Tier MDPs for simplicity, and generalize to $k$-Tier MDPs in Section 6. In a 3-Tier MDP, we will call the 3 tiers obstacles ($S_{1}$), background ($S_{2}$), and goals ($S_{3}$), in order of increasing desirability. States in $S_{1}$ and $S_{3}$ are absorbing. We define $o_{t}$ to be the probability of being in $S_{1}$ at timestep $t$, and $g_{t}$ that of $S_{3}$.
@@
Given two policies $\pi^{A}$ and $\pi^{B}$, we say $\pi^{A}$ dominates $\pi^{B}$ when both of these inequalities hold (and not both being strictly equal at all times):
$\sum_{i=0}^{t}o_{i}^{A}\leq\sum_{i=0}^{t}o_{i}^{B},\quad\forall t=0,1,2,...,\infty,$
$\sum_{i=0}^{t}g_{i}^{A}\geq\sum_{i=0}^{t}g_{i}^{B},\quad\forall t=0,1,2,...,\infty.$
@@
In words, one policy dominates another if it gets to the goal faster, while delaying encountering obstacles longer. The set of policies that are not dominated by any other policy is the set of Pareto-optimal policies. Because there is a finite number of policies and domination is transitive, the set of Pareto-optimal policies is non-empty. We provide a visualization of Pareto-optimal policies in Appendix C.
@@
Pareto-optimal policies are interesting to consider for two main reasons. First, Pareto-optimal behavior always exists, even when policies that achieve other reasonable things do not. Secondly, Pareto-optimality resolves the preference problem by defining a strict partial ordering over the entire policy space. Although the policies on the Pareto frontier are incomparable among themselves, they are all better than the set of Pareto-dominated policies. We simply deem the set of Pareto-optimal policies to be the desirable behavior, and all others undesirable. Next, we show how to design rewards that guarantee Pareto-optimal policies.
@@ 5 Tiered Reward
In this section, we seek a sufficient condition on the reward function so that optimizing expected discounted reward will always result in a Pareto-optimal policy with respect to our preference relation.
@@
Table 1: Three example reward functions of Russell/Norvig grid world.
In fact, many of the reward functions that satisfy $r_{lava}<r_{background}<r_{goal}$ are not Pareto-optimal. Out of $1000$ such rewards that we sampled randomly, 90.5% were Pareto-dominated (See Figure 13 in Appendix). Next, we present a simple rule that is sufficient to guarantee environment-independent Pareto-optimal reward functions in 3-Tier MDPs.
@@
6 Generalizing to $k$ Tiers
In Sections 4 and 5, we limited the discussion to $3$-Tier MDPs. But MDPs with more than $3$ tiers can usefully model important problems such as those with well-defined subgoal states. Specifically, each subgoal region could be its own tier, instead of being grouped into one big background tier. Even though these problems could still be solved as a $3$-Tier MDP, more knowledge about the environment could help design better reward functions and accelerate learning. So, in this section, we consider the reward-design problem in Tier MDPs with more than three tiers.
@@ 6 Generalizing to $k$ Tiers
Notice that the Tiered Reward in the $k$ tiers case uses a stricter condition than that of $3$ tiers. First, all reward values are non-positive. This is a sufficient but not necessary condition to guarantee Pareto-optimality. We enforce this constraint not only because of mathematical convenience (used later in Equation 1), but also because step-wise penalty has been proved to support faster learning (Koenig and Simmons 1993). Specifically, with a zero-initialized value function, step penalties create an incentive for the agent to try state–action pairs it has never experienced before, resulting in rapid exploration. Secondly, the reward values of higher tiers are exponentially greater than the lower ones. For adjacent tiers $i$ and $i+1$, The reward values always satisfy $r_{i}<\frac{1}{1-\gamma}r_{i+1}<0$. One such reward is visualized in Figure 2.
This definition can be understood as a generalization of the 3-Tiered Reward. When the agent resides within tier $i\in\{2,3,...,k-1\}$, the $k$ tiers could be partitioned into $3$ groups to construct a 3-Tier MDP. In particular, $S_{1}$ will include tiers $1$ through $i-1$, $S_{2}$ is just tier $i$, and $S_{3}$ is tiers $i+1$ to $k$. Note that we can generalize Theorem 5.3 to allow states in $S_{1}$ and $S_{3}$ to have any reward values as long as they satisfy the inequality in Definition 5.2 for a fixed reward value in $S_{2}$. Namely, denote $r_{low}=\max\{r_{1},...,r_{i-1}\}$ and $r_{high}=\min\{r_{i+1},...,r_{k}\}$, and as a $k$-Tiered Reward they satisfy
$r_{low}<(\frac{1}{1-\gamma})r_{i}<(\frac{1}{1-\gamma})^{2}r_{high}$
And since $\gamma\in(0,1)$,
$r_{low}<(\frac{1}{1-\gamma})r_{i}<(\frac{1}{1-\gamma})^{2}r_{high}\leq r_{high}.$ (1)
That is, $(r_{low},r_{i},r_{high})$ is a Tiered Reward function in the $3$-Tier MDP with tiers $S_{1}$, $S_{2}$, and $S_{3}$, and therefore induces Pareto-optimal policies (Theorem 5.3). So, at tier $i$, the policy that optimizes the $k$-Tiered Reward will push agents to higher tiers as fast as possible and avoid lower tiers, as if they were goals and obstacles, respectively. In the special case that the agent resides within tier $i=1$, the constraint from Definition 6.1 will treat tiers $2$ through $k$ as if they are all goals, pushing the agent towards them. In the case that $i=k$, the agent is already in the “goal tier”. So overall, $k$-Tiered Reward will induce in a ratchet-like policy—go to the higher tiers as fast as possible while not fall back to the lower tiers—that makes learning fast. In fact, it has been shown that a similar increasing reward profile (Sowerby, Zhou, and Littman 2022) leads to fast learning. Okudo and Yamada (2021) and Zhai et al. (2022) have also shown that intermediate rewards can accelerate learning and provably improve sample efficiency in goal–reaching tasks.
@@
[…]
$\sum_{i=0}^{t}p_{i}^{d}\geq\sum_{i=0}^{t}{p_{i}^{*}}^{d},\quad\forall d=[2..k],\forall t=0,1,2,...,\infty.$

### v3

# Tiered Reward: Designing Rewards for Specification and Fast Learning of Desired Behavior

@@ Abstract
Reinforcement-learning agents seek to maximize a reward signal through environmental interactions. As humans, our job in the learning process is to design reward functions to express desired behavior and enable the agent to learn such behavior swiftly. However, designing good reward functions to induce the desired behavior is generally hard, let alone the question of which rewards make learning fast. In this work, we introduce a family of a reward structures we call Tiered Reward that addresses both of these questions. We consider the reward-design problem in tasks formulated as reaching desirable states and avoiding undesirable states. To start, we propose a strict partial ordering of the policy space to resolve trade-offs in behavior preference. We prefer policies that reach the good states faster and with higher probability while avoiding the bad states longer. Next, we introduce Tiered Reward, a class of environment-independent reward functions and show it is guaranteed to induce policies that are Pareto-optimal according to our preference relation. Finally, we demonstrate that Tiered Reward leads to fast learning with multiple tabular and deep reinforcement-learning algorithms.
@@ 1 Introduction
Reinforcement learning (Sutton & Barto, 1998) (RL) is concerned with the problem of learning to behave to maximize a reward signal. In biological systems, this reward signal is considered to be the organism’s motivational system, using pain and pleasure to modulate behavior (Porreca & Navratilova, 2017; Leknes & Tracey, 2010; Navratilova & Porreca, 2014). In engineered systems, however, rewards must be selected by the system designer (Nagpal et al., 2020; Dewey, 2014). We view rewards as a kind of programming language—a specification of the agent’s target behavior (Littman et al., 2017; Zhou & Li, 2022). As arbiters of behavior correctness in the learning process, humans bear the responsibility of authoring this program. This is referred to as the reward-design problem (Dewey, 2014; Devidze et al., 2021; Sorg et al., 2010; Sowerby et al., 2022): given a set of desired behavior, what kind of reward functions would efficiently express these behavior? In this paper, we look at rewards that can correctly and efficiently express desirable states (goals and subgoals) and undesirable states (obstacles).
@@
There are two essential steps in designing reward functions. First, one must decide what kind of behavior is desirable and should be conveyed. Then, there’s the choice of reward function that induces such behavior. The first step is hard because there is no universal preference over behavior and having to explicitly write down all possible trade-offs is challenging. Even if the reward designer has a way of expressing preferences for all possible exchanges, it can be difficult, or even impossible, to design a reward function that captures them without prior knowledge of the environment.
@@
First, let us consider the question of how to specify preference over a set of policies (Roy et al., 2021). In the goal–obstacle class of tasks we consider, preferences over policies are simple in deterministic environments. We imagine all states are either goal states, obstacle states, or neither (background states), and all goals and obstacles are absorbing. Preferences in deterministic environments form a total order: reaching goals is better than reaching obstacles; if the policy reaches goals, faster is better; if the policy reaches obstacles, avoiding them longer is better. However, even in this simple setting, providing such precise trade-offs is difficult in stochastic environments. Is it better for an agent to increase the chance of getting to the goal by 5% if it also incurs 8% higher probability of hitting an obstacle? Is it better to increase by 50% the probability of getting to a goal if the expected time of getting there also increases by 20%? Preferences are less clear in a stochastic setting because there can be trade-offs between different outcomes and their probabilities. However, some comparisons are arguably clear cut. Informally, if one policy induces uniformly better outcomes than another—being more likely to reach a goal and doing so faster, being less likely to reach an obstacle and getting there more slowly—we prefer such a policy. If the policies can’t be directly compared, we propose to be indifferent between them. Thus, we replace the standard reinforcement-learning notion of optimality with Pareto-optimality (Mornati, 2013) which is commonly adopted in multiobjective RL (Vamplew et al., 2011; Van Moffaert & Nowé, 2014); we seek a policy that is either preferred or incomparable to every other policy.
@@
Even after resolving the issue of specifying behavior preference, policies are hard to express through reward functions in general (Zhang et al., 2009; Amodei & Clark, 2016), and some are even impossible to convey with a Markov (state–action-based) reward (Abel et al., 2021). Even when policies are expressible, designing bad reward functions can lead to undesirable or dangerous actions (Amodei & Clark, 2016), easy reward hacking (Amodei et al., 2016; Skalse et al., 2022), and more. We seek to design good reward functions, which can be characterized by many properties, such as interpretability and learning speed (Devidze et al., 2021). But the most important property a reward function must have is to guarantee the adoption of a desired policy. As we will show later in Section 5, even intuitively correct reward designs can lead to suboptimal policies. To hedge against this, we introduce a tiered reward structure that is guaranteed to induce Pareto-optimal policies. Intuitively, we partition the state space into several tiers, or goodness levels. States in the same tier are associated with the same reward, while states in a more desirable tier are associated with an exponentially higher reward. We prove that these tiered reward structures, with the proper constraints between reward values, induce Pareto-optimal behavior and empirically show that they can lead to fast learning.
@@
In this paper, we propose Tiered Reward as a way to design reward functions that can correctly and swiftly express desired behavior. Our contribution is threefold: First, we define a preference over the entire policy space via a strict partial ordering on outcomes using the notion of Pareto-optimality. This addresses the question of behavior preference in stochastic environments. Then, we introduce a class of environment-independent tiered reward structure that provably induce Pareto-optimal policies with respect to this preference ordering. Finally, we demonstrate these tiered reward functions can lead to fast learning in both tabular and deep RL settings and is invariant to the choice of RL algorithms.
@@ 2 Related Work
Specifying behavior through rewards:
Preference-based RL methods (Wirth et al., 2017; Brown et al., 2019; Liu et al., 2022) learn a reward function based on a dataset of preferences over trajectories. But, as we have argued, preferences over trajectory probabilities can be very difficult to specify. In addition, our reward scheme relieves the need for environment-specific preference datasets created by human experts. Multi-objective RL (Vamplew et al., 2011; Toro Icarte et al., 2018; Hayes et al., 2022) allows for different tasks to be specified through a set of reward functions. Our work proceeds in the orthogonal direction by designing a single reward function to trade off among multiple behaviors, instead of incentivizing all of them. Reward machines (Icarte et al., 2018; 2022) are finite state machines that compose reward functions and allow different rewards to be delivered dependent on the agent’s trajectory. They reveal the structure of the reward function to the RL agent to support decomposition of complex tasks. Our focus on how to provide incentives for specific outcomes is complementary and the two approaches can be used in concert. Temporal logic based languages (Littman et al., 2017; Camacho et al., 2017; Li et al., 2017; Camacho et al., 2019) have been used to specify behavior. Though these methods can be more expressive, they often lead to intractable planning and learning problems due to state-space explosion issues (Wongpiromsarn et al., 2010). We offer a different expressibility–tractability tradeoff.
Reward Design for fast learning:
Mataric (1994) proposed to accelerate learning by incorporating domain knowledge and using a progress estimator, but does not provide a principled method of designing a reward function. Sowerby et al. (2022) showed that reward functions that maximize the action gap given a measure of horizon length lead to fast learning. However, designing such reward functions requires solving an optimization problem with detailed knowledge of the environment, making this approach impractical. Similarly, Devidze et al. (2021) formulated the reward-design problem as an optimization problem to maximize informativeness and sparsity. However, their method requires solving the MDP with ground truth transition dynamics and a reference reward function, which is often not available in practice.
@@ 3 Problem Setting
We are interested in the reward-design problem. In the common RL framework, tasks are specified by the reward function and the agent’s objective is to maximize cumulative rewards. We take an alternative perspective: We specify tasks by prescribing a set of desirable policies and seek reward functions such that maximizing the reward will lead to the desirable policies. Formally, given a set of desirable (Pareto-optimal) policies $\Pi$, the reward-design problem is to create a reward function $R:S\rightarrow R$ such that the optimal policy $\pi_{R}^{*}\in\Pi$.
We imagine the state space $S$ as exhibiting a tiered structure, where higher tiers are more desirable than lower tiers, and states within the same tier are equally desirable. Formally, we define:
@@
(a) Russell/Norvig grid world. Objective is to reach the goal (green) without first visiting lava (orange) ($\gamma=0.9$).
(b) A Tiered Reward in a grid world with $6$ tiers. Start state in bottom left, goal state in top right. Darker colors correspond to more negative reward values.
Figure 1:
@@
As an example, the grid world from Russell & Norvig (2010), as illustrated in Figure 1(a), could be formulated as a 3-Tier MDP—the goal state is one tier ($S_{3}$), the lava state one tier ($S_{1}$), and all other states reside in the background tier ($S_{2}$). It is important to note that we put no constraints on how many states could be in each tier, nor how many tiers there can be. Therefore, the framework has a high degree of generality: any finite MDP with reward defined on states could be formulated as a Tier MDP by placing states with the same reward in the same tier. However, the Tier MDP is most useful when there are clear good and bad states in the state space, such as when there are goal and obstacle states, or even states of intermediate desirability such as subgoal states. In the following sections, we will show how to perform reward design in Tier MDPs.
4 Policy Ordering with Pareto-optimality
@@
A policy induces a probability distribution over an infinite set of outcomes (specifically the probability of reaching each of the states after $t$ steps, for all $t$). In goal–obstacle tasks, policies can be characterized by statistics such as probability of reaching the goal and probability of avoiding the obstacle for each possible horizon length. Using these statistics, we will show below how the entire policy space can form a strict partial ordering to specify which policies are preferable.
@@
We define $o_{t}$ to be the probability of being in $S_{1}$ at timestep $t$, and $g_{t}$ that of $S_{3}$. Given two policies $\pi^{A}$ and $\pi^{B}$, we say $\pi^{A}$ dominates $\pi^{B}$ when both of these inequalities hold (and not both being strictly equal at all times):
$\sum_{i=0}^{t}o_{i}^{A}\leq\sum_{i=0}^{t}o_{i}^{B},\quad\sum_{i=0}^{t}g_{i}^{A}\geq\sum_{i=0}^{t}g_{i}^{B},\quad\quad\forall t=0,1,2,...,\infty.$
@@
In words, one policy dominates another if it gets to the goal faster, while delaying encountering obstacles longer. The set of policies that are not dominated by any other policy is the set of Pareto-optimal policies. Because there is a finite number of policies and domination is transitive, the set of Pareto-optimal policies is non-empty. See a visualization of Pareto-optimal policies in Appendix C.
@@
Pareto-optimal policies are interesting to consider for two main reasons. First, Pareto-optimal behavior always exists, even when policies that achieve other reasonable things do not. Secondly, Pareto-optimality addresses the preference problem by defining a strict partial ordering over the entire policy space. Although the policies on the Pareto frontier are incomparable among themselves, they are all better than the set of Pareto-dominated policies. We simply deem the set of Pareto-optimal policies to be the desirable behavior, and all others undesirable. Next, we show how to design rewards that guarantee Pareto-optimal policies.
@@ 5 Tiered Reward
In this section, we seek a sufficient condition on the reward function so that optimizing expected discounted reward will result in a Pareto-optimal policy with respect to our preference relation.
@@
Table 1: Three intuitive reward functions of Russell/Norvig grid world. $B$ is Pareto-dominated.
In fact, many of the reward functions that satisfy $r_{lava}<r_{background}<r_{goal}$ are not Pareto-optimal. Out of $1000$ such rewards that we sampled randomly, 90.5% were Pareto-dominated (See Figure 11 in Appendix). Next, we present a simple rule that is sufficient to guarantee environment-independent Pareto-optimal reward functions.
5.1 The $3$ Tiers Case
To facilitate understanding, we limit the problem space to 3-Tier MDPs for now, and generalize to $k$-Tier MDPs in Section 5.2. In a 3-Tier MDP, we will call the 3 tiers obstacles ($S_{1}$), background ($S_{2}$), and goals ($S_{3}$), in order of increasing desirability. States in $S_{1}$ and $S_{3}$ are absorbing.
@@ 5.1 The $3$ Tiers Case
5.2 Generalizing to $k$ Tiers
MDPs with more than $3$ tiers can usefully model important problems such as those with well-defined subgoal states. Specifically, each subgoal region could be its own tier, instead of being grouped into one big background tier. Even though these problems could still be solved as a $3$-Tier MDP, more knowledge about the environment could help design better reward functions and accelerate learning.
@@ 5.2 Generalizing to $k$ Tiers
One such reward is visualized in Figure 1(b). Notice that the $k$-Tiered Reward (Definition 5.4) uses a stricter condition than that of $3$-Tiered Reward (Definition 5.2). In fact, Definition 5.4 is a sufficient condition for Definition 5.2 because $3$ tiers is a special case with only one non-absorbing tier. One can also design a stricter $3$-Tiered Reward with Definition 5.4 with $k=3$. Definition 5.4 is stricter in that: first, all reward values are non-positive. This is a sufficient but not necessary condition to guarantee Pareto-optimality. We enforce this constraint not only to make it a sufficient condition, but also because step-wise penalty has been proved to support faster learning (Koenig & Simmons, 1993; Sowerby et al., 2022). Specifically, with a zero-initialized value function, step penalties create an incentive for the agent to try state–action pairs it has never experienced before, resulting in rapid exploration. Secondly, the reward values of higher tiers are exponentially greater than the lower ones. For adjacent tiers $i$ and $i+1$, The reward values always satisfy $r_{i}<\frac{1}{1-\gamma}r_{i+1}<0$.
This definition can be understood as a generalization of the 3-Tiered Reward. When the agent resides within tier $i\in\{2,3,...,k-1\}$, the $k$ tiers could be partitioned into $3$ groups to construct a 3-Tier MDP. In particular, $S_{1}$ will include tiers $1$ through $i-1$, $S_{2}$ is just tier $i$, and $S_{3}$ is tiers $i+1$ to $k$. Note that we can generalize Theorem 5.3 to allow states in $S_{1}$ and $S_{3}$ to have any reward values as long as they satisfy the inequality in Definition 5.2 for a fixed reward value in $S_{2}$. Namely, denote $r_{low}=\max\{r_{1},...,r_{i-1}\}$ and $r_{high}=\min\{r_{i+1},...,r_{k}\}$, and as a $k$-Tiered Reward they satisfy $r_{low}<(\frac{1}{1-\gamma})r_{i}<(\frac{1}{1-\gamma})^{2}r_{high}$ And since $\gamma\in(0,1)$, $r_{low}<(\frac{1}{1-\gamma})r_{i}<(\frac{1}{1-\gamma})^{2}r_{high}\leq r_{high}.$ That is, $(r_{low},r_{i},r_{high})$ is a Tiered Reward function in the $3$-Tier MDP with tiers $S_{1}$, $S_{2}$, and $S_{3}$, and therefore induces Pareto-optimal policies (Theorem 5.3). So, at tier $i$, the policy that optimizes the $k$-Tiered Reward will push agents to higher tiers as fast as possible and avoid lower tiers, as if they were goals and obstacles, respectively. In the special case that the agent resides within tier $i=1$, the constraint from Definition 5.4 will treat tiers $2$ through $k$ as if they are all goals, pushing the agent towards them. In the case that $i=k$, the agent is already in the “goal tier". So overall, $k$-Tiered Reward will induce in a ratchet-like policy—go to the higher tiers as fast as possible while not fall back to the lower tiers—that makes learning fast. In fact, it has been shown that a similar increasing reward profile (Sowerby et al., 2022) leads to fast learning. Okudo & Yamada (2021) and Zhai et al. (2022) have also shown that intermediate rewards can accelerate learning and provably improve sample efficiency in goal–reaching tasks.
@@
$\sum_{i=0}^{t}p_{i}^{1}\leq\sum_{i=0}^{t}{p_{i}^{*}}^{1},\quad\forall t=0,1,2,...,\infty\quad\quad\sum_{i=0}^{t}p_{i}^{d}\geq\sum_{i=0}^{t}{p_{i}^{*}}^{d},\quad\forall d=[2..k],\forall t=0,1,2,...,\infty.$